# Reparameterized LLM Training via Orthogonal Equivalence Transformation

### v2

@@ 2 From Energy-preserving Training to Spectrum-preserving Training
POET preserves hyperspherical energy under isotropic Gaussian initialization. [47] shows that weight matrices that are initialized by zero-mean isotropic Gaussian distribution (e.g.[16]) are guaranteed to have small hyperspherical energy. Because zero-mean isotropic Gaussian is invariant under orthogonal transformation, $\bm{R}\bm{W}\bm{P}$ also has small hyperspherical energy (see Appendix B).
@@ 4 Discussions and Intriguing Insights
Guarantees of weight spectrum. For POET with standard and normalized Gaussian initializations, we prove in Appendix C that the largest and smallest singular values of weights can be bounded.
@@ 5.2 Ablation Studies and Empirical Analyses
Initialization schemes. We empirically compare different random initialization schemes for POET, including two commonly used ones (standard Gaussian, Xavier [16]) and two proposed ones (uniform spectrum, normalized Gaussian). Specifically, we use fully stochastic POET with $b$=1/2 to train Llama-60M on 30B tokens and report the validation perplexity in Table 3. Results shows that the normalized initialization will lead to the best final performance, and we stick to it as a default choice. Interestingly, uniform spectrum initialization performs poorly. This suggests a trade-off between preserving good weight spectral properties and achieving strong expressiveness. it may limit its expressiveness. Finding the optimal singular value structure for weights remains an important open problem.

### v3

@@ 2 From Energy-preserving Training to Spectrum-preserving Training
POET preserves hyperspherical energy under isotropic Gaussian initialization. [48] shows that weight matrices that are initialized by zero-mean isotropic Gaussian distribution (e.g.[17]) are guaranteed to have small hyperspherical energy. Because zero-mean isotropic Gaussian is invariant under orthogonal transformation, $\bm{R}\bm{W}\bm{P}$ also has small energy (see Section 4 and Appendix B).
@@ 4 Discussions and Intriguing Insights
Guarantees of weight spectrum. For POET with standard and normalized Gaussian initializations, we have proved in Appendix C that the largest and smallest singular values of weights can be bounded. For normalized Gaussian, the bound is only dependent on the row-to-column ratio of the weight matrix. For standard Gaussian, the bound has an extra dependence on the neuron dimension.
@@
Figure 6: Dynamics comparison of average singular value entropy (singular value diversity) between direct training (AdamW, Muon) and POET.
Dynamics of singular spectrum. Inspired by [46], we conduct a spectral analysis by comparing the singular spectrum among AdamW, Muon and POET. Following [1, 68], we compute SVD entropy of the trained Llama-60M model at different iteration. Specifically, the SVD entropy, defined as $H(\bm{\sigma})=\frac{1}{\log n}\sum_{i}\frac{\sigma_{i}^{2}}{\sum_{j}\sigma_{j}^{2}}\log\frac{\sigma_{i}^{2}}{\sum_{j}\sigma_{j}^{2}}$ measures the diversity of singular values; higher entropy indicates a more uniform and diverse spectrum. [46] attributes the favorable performance of Muon over AdamW to the more diverse spectrum of weight matrices updates. As shown in Figure 6, POET consistently maintains high spectral diversity throughout training, owing to its orthogonal equivalence transformation. Therefore, POET can better explore diverse optimization directions. Consistent with the observation in [46], we also find that Muon yields more diverse spectral updates than AdamW.
POET minimizes hyperspherical energy in expectation. While POET effectively generalizes energy-preserving training to spectrum-preserving training, it still ensures low hyperspherical energy when initialized with a zero-mean isotropic Gaussian distribution. This is actually a significant property, as POET retains the generalization benefits of minimal hyperspherical energy [47, 49, 48]. This property comes from the invariance of zero-mean isotropic Gaussian to orthogonal transformation. [48] shows that for a weight matrix $\bm{W}$ that is initialized by zero-mean isotropic Gaussian, its neurons, after being normalized, are uniformly distributed on the unit hypersphere. This property provably leads to a small hyperspherical energy. Since zero-mean isotropic Gaussian initialization is invariant to orthogonal transformation, $\text{OET}(\bm{W};\bm{R},\bm{P})=\bm{R}\bm{W}\bm{P}$ does not change the distribution of $\bm{W}$, i.e., $\bm{R}\bm{W}\bm{P}=_{d}\bm{W}$. We give a detailed derivation in Appendix B. Therefore, the transformed neurons of the equivalent weight matrix $\bm{R}\bm{W}\bm{P}$ are also uniformly distributed over the unit hypersphere after being normalized. This validates that POET can provably preserve a small energy.
Figure 7: Hyperspherical energy comparison between AdamW, Muon and POET.
The property that POET preserves singular spectrum while retaining minimal hyperspherical energy in expectation well justifies why POET yields stable training and good generalization performance. Such a property only holds true when the weight initialization follows zero-mean isotropic Gaussian, which is perfectly satisfied by current weight initialization. It may also justify why zero-mean isotropic Gaussian initialization (e.g., standard Gaussian, normalized Gaussian, Xavier [17]) works better than the uniform spectrum initialization in our experiments. To further validate that POET maintains a small hyperspherical energy, we plot the hyperspherical energy of the Llama-60M model at different iterations in Figure 7. We compare the sum of the hyperspherical energy of all the layers trained by AdamW, Muon or POET. The results again verify POET’s energy-minimizing property. Slightly different from previous orthogonal training [48], POET can not preserve exactly the same hyperspherical energy throughout the training, but it can provably minimize hyperspherical energy in an expectation sense. Interestingly, we find that Muon can also minimize hyperspherical energy more effectively than AdamW, which partially justifies why Muon may lead to better performance.
@@ 5.2 Ablation Studies and Empirical Analyses
Initialization schemes. We empirically compare different random initialization schemes for POET, including two commonly used ones (standard Gaussian, Xavier [17]) and two proposed ones (uniform spectrum, normalized Gaussian). Specifically, we use fully stochastic POET with $b$=1/2 to train Llama-60M on 30B tokens and report the validation perplexity in Table 3. Results show that the normalized initialization will lead to the best final performance, and we stick to it as a default choice. Interestingly, uniform spectrum initialization performs poorly. This suggests a trade-off between preserving good weight spectral properties and achieving strong expressiveness. it may limit its expressiveness. Finding the optimal singular value structure for weights remains an important open problem.